OpenStudy (anonymous):

Can someone help me with this one? Find the indicated outputs for f(x)=3X2-2X F(0)= F(-1)= F(2)=

OpenStudy (anonymous):

0, 5, 8

OpenStudy (anonymous):

replace x by the numbers given. so for example \[f(-1)=3(-1)^2-2\times (-1)=3+2=5\]

OpenStudy (anonymous):

use parentheses and be careful with the signs